Grade 12th passMechanics

A ball is thrown vertically upwards with an initial velocity of 36m/s.After 2sec another ball is thrown vertically upwards, determine the initial velocity of the ball when it crosses the first ball at 30m ?

Ashim Kumar
9 Years agoGrade 12th pass

1 Answer

Aditya kulkarni
9 Years ago
For a ball AU- 36m/sHence By using kinemetical equationS=ut-1/2gt^230= 36t -5t^2........(taking g=10)By factorsing given equationYou will get 6.2 and 1As the ball 2 is thrown 2sec. Later...time taken by him will be(T+2) hence putting the values of t in the equation30=u(t+2) -1/2g(t+2)^2By putting t=6.2 u will get u=41m/sNd by putting t=1 u will get u= 25m/sAs it is not valid! Because it should be greater than 36 as it is thrown later